OpenStudy (anonymous):

For all rational numbers r and s if s is nonzero then 2r/5s is rational. Still confused

OpenStudy (skullpatrol):

What is the definition of a rational number?

OpenStudy (skullpatrol):

In math you use definitions to clarify what something means.

OpenStudy (skullpatrol):

@jamjam1988 are you still there?

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (ikram002p):

the rational number is the number that u can write it as the form of a/b r is rational r=a/b s is rational s=c/d 2r/5s=2(a/b)/5(c/d) \[\frac{ 2r }{ 5s }=\frac{ 2\frac{ a }{ b } }{ 5\frac{ c }{d } }\] \[=\frac{ 2a }{ b }\times \frac{ d }{ 5c }\] =\[=\frac{ 2ad }{5bc }\] assume 2ad=k 5bc=h \[\frac{ 2r }{5s }=\frac{ k }{h }\] #

OpenStudy (anonymous):

yes im here would not let me reply